ICM 1998, Berlin, Aug. 18–27 Abstracts of Plenary and Invited Lectures

Section:

16. Applications

Section 16 a: Applications of Mathematics in Other Sections. 1991 MS Classification: Avellaneda, Marco, Courant Institute, New York University, USA Inverse Problems in Quantitative Financial Modeling (see section 12)

ICM 1998 Berlin · Plenary and Invited Lectures